Solve the simultaneous differential equation
dx/x^2-y^2-z^2=dy/2xy=dz/2xz
1
Expert's answer
2020-11-29T19:16:59-0500

Given differential equation is, "\frac{dx}{x^2-y^2-z^2}=\frac{dy}{2xy}=\frac{dz}{2xz}"



Taking last two terms, "\frac{dy}{2xy}=\frac{dz}{2xz} \implies \frac{dy}{y}=\frac{dz}{z}"

Integrating both sides, we get,

"y=cz" (1)


Taking first and last term,

"\frac{dx}{x^2-y^2-z^2}=\frac{dz}{2xz}"


"\frac{dx}{x^2-c^2z^2-z^2}=\frac{dz}{2xz}"


"\frac{dx}{dz} = \frac{x^2-z^2(1+c^2)}{2xz}"


Putting x=vz, then, "\frac{dx}{dz} = v+z\frac{dv}{dz}"


Then equation will be, "v+z\frac{dv}{dz} = \frac{v^2x^2-z^2(1+c^2)}{2vz^2} = \frac{v^2-(1+c^2)}{2v}"

"z\frac{dv}{dz} = -\frac{v^2+(1+c^2)}{2v}"


"\frac{2vdv}{v^2+(1+c^2)} = -\frac{dz}{z}"

Integrating the above equation, "log|v^2+(1+c^2)|=-log|z|+log|c_1| \implies v^2+(1+c^2)=\frac{c_1}{z}"

Putting value of v, "\frac{x^2}{z^2}= \frac{c_1}{z} -(1+c^2)"

"x^2 = c_1z - z^2(1+c^2)"


Solving quadratic equation for z, we get,


"z=\frac{c_1\pm \sqrt{(c_1^2-4x^2(1+c^2)})}{ 2(1+c^2)}" (2)


Equation (1) & (2) are solutions.